A rectangular block on material is subjected to a tensile stress of 100 N/mm2 on one plane and

a tensile stress of 50 N/mm2

on a plane at right angles, together with shear stress of 60N/mm2

on the

faces. Find:

a) The direction of principal planes

b) The magnitude of principal stresses and

c) Magnitude of greatest shear stress.


1
Expert's answer
2021-02-28T07:34:09-0500

Stress in x direction:

"\sigma_x=100\ N/mm^2"

Stress in y direction:

"\sigma_y=50\ N/mm^2"

Shear stress:

"\tau_{xy}=60\ N/mm^2"


Principal stresses:

"\sigma_{1,2}=\frac{\sigma_x+\sigma_y}{2}\pm\sqrt{(\frac{\sigma_x-\sigma_y}{2})^2+\tau_{xy}^2}"

"\sigma_{1,2}=\frac{100+50}{2}\pm\sqrt{(\frac{100-50}{2})^2+60^2}"

"\sigma_1=140\ N/mm^2"

"\sigma_2=10\ N/mm^2"


Direction of principal plane:

"tan2\theta_p=-\frac{\sigma_x-\sigma_y}{2\tau_{xy}}"

"tan2\theta_p=-\frac{100-50}{2\cdot60}=-5/12"

"\theta_p=-11.31\degree"


The greatest shear stress:

"\tau_{max}=|\sigma_1-\sigma_2|/2"

"\tau_{max}=(140-10)/2=65\ N/mm^2"